OpenStudy (steparti1):

Chloe lives in an area where she runs the air conditioner during the 4 hottest months of the year. The air conditioner runs for 7 hours per day at a cost of 21 cents per hour. How much does she spend per year on utility bills for the air conditioner?(Assume there are 30 days in each month) a. $705.60 b. $588.00 c. $216.09 d. $176.40

OpenStudy (anonymous):

4 months, each month is 30 days. 30*4=120 She has the AC on for 120 days. Each day, she has the AC on for 7 hours. 7*120=840 total hours Chloe's AC is on in a year. AC costs 21 cents per hour. Remember, 21 cents would be written as $0.21, not $21! 840*0.21=$176.40 spent on AC per year